OpenStudy (anonymous):

identify the terms and the coefficients of the expression 18-2x-4y. since the expression involves________ rather than addition, rewright the expression as the____ of the terms: 18-2x-4y= ____________________. so the terms of the expression are ___________ and the coefficients are _____.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

#jaycindagraun

OpenStudy (anonymous):

@jaycindagraun

OpenStudy (anonymous):

Did you find it out ?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

A subtraction can be written as an addition of the opposite. In general, a - b = a + (-b) Example: 5x - 3y = 5x + (-3y) A polynomial is a SUM of terms. If a polynomial shows the terms being subtracted, rewrite the subtractions as additions of the opposite, and that shows the terms. Example: What are the terms of the following polynomial? \(2x^3 - 3x^2 + 4x - 7\) Solution. We rewrite the polynomial showing the sum of the terms: \(2x^3 + (-3x^2) + 4x + (-7)\) The terms are: \(2x^3,~-3x^2, ~4x, ~-7\) What this question is doing is to make sure you understand that each term includes the minus sign to its left.

OpenStudy (anonymous):

but what goes in the blanks